OpenStudy (anonymous):

3x + 3y=10 -9x - 9y=-30 solve please!!!!!!!!!!!1

OpenStudy (anonymous):

do you know how to work simultaneous equations?

OpenStudy (anonymous):

not really, can you show me with different numbers?

OpenStudy (anonymous):

no, these will do just fine, make x or y in either one of the equations the subject of the formula

OpenStudy (paounn):

Second equation is the first multiplied by -3. They are in fact the same equation, and the system admits infinte^1 solutions. Fix either x or y as t, and solve on the other.

OpenStudy (anonymous):

so i will subtract 3x from both sides on the first equation?

OpenStudy (anonymous):

yea u r going right, go brave and show me ur final answer!

OpenStudy (anonymous):

3y=10+3x?

OpenStudy (anonymous):

oops 3y=10-3x

OpenStudy (anonymous):

gud! now u want it to be just y on one side so now divide both sides by 3! and give meh ur answer!

OpenStudy (anonymous):

-3x?

OpenStudy (anonymous):

-3.3x?

OpenStudy (anonymous):

u'll get y = 10/3 - x :)

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

yes is that the answer to both of the problems?

OpenStudy (anonymous):

nope, now u have to substitute it into the other equation

OpenStudy (anonymous):

substitute what?

OpenStudy (anonymous):

i will give a medal to anyone who responds with explanations